OpenStudy (anonymous):

write the equation of a parabola with the shape as f(x)= -2xsquared but with vertex (-3,2)

OpenStudy (anonymous):

put y instead of f(x), the equation is \[y=-2x^2\] with vertex (-3,2) and the same shape, (just change the vertex from (0,0) to (-3,2) ): \[y-2=-2(x-(-3))^2 \implies y-2=-2(x+3)^2\]

OpenStudy (anonymous):

A ball is thrown upward from a 100-foot tall building with an initial velocity of 14 feet per second. Its height s(t) is given by the function s(t) = -16t squared + 14t + 100. Find the interval of time for which the ball is greater than 103 feet.

OpenStudy (anonymous):

A whole number increased by its square is 4 times itself. Find the number.

OpenStudy (anonymous):

A square picture frame will hold a poster with an area of 196 square inches. Find the dimensions of the frame.

OpenStudy (anonymous):

you should write each question separated in a different page!

OpenStudy (anonymous):

ok